OpenStudy (anonymous):

4x²-4 = 4 Solve quadratic function with square roots

OpenStudy (xapproachesinfinity):

Hint add 4 to both sides

OpenStudy (anonymous):

Then I get 4x²=8 Do I divide by 4 now?

OpenStudy (xapproachesinfinity):

Exactly!

OpenStudy (anonymous):

So then x²=2 How do I square root that? Or is there no way

OpenStudy (xapproachesinfinity):

apply square root to both sides

OpenStudy (xapproachesinfinity):

\[\sqrt{x^2}=\sqrt{2}\]

OpenStudy (anonymous):

And having a decimal as an answer is fine?

OpenStudy (xapproachesinfinity):

no leave it like that so what are the answers? they two answers

OpenStudy (xapproachesinfinity):

there are two answers

OpenStudy (anonymous):

x=-2 x=2 ?

OpenStudy (xapproachesinfinity):

not quite they are \[x=\sqrt{2} , ~ x=-\sqrt{2}\]

OpenStudy (xapproachesinfinity):

i don't why you destroyed that square root you needed to simplify the left side \[\sqrt{x^2}\]